### v3

# TOP-Nav: Legged Navigation Integrating Terrain, Obstacle and Proprioception Estimation

Abstract
Legged navigation is typically examined within open-world, off-road, and challenging environments. In these scenarios, estimating external disturbances requires a complex synthesis of multi-modal information. This underlines a major limitation in existing works that primarily focus on avoiding obstacles. In this work, we propose TOP-Nav, a novel legged navigation framework that integrates a comprehensive path planner with Terrain awareness, Obstacle avoidance and close-loop Proprioception. TOP-Nav underscores the synergies between vision and proprioception in both path and motion planning. Within the path planner, we present a terrain estimator that enables the robot to select waypoints on terrains with higher traversability while effectively avoiding obstacles. In the motion planning level, we construct a proprioception advisor from the learning-based locomotion controller to provide motion evaluations for the path planner. Based on the close-loop motion feedback, we offer online corrections for the vision-based terrain and obstacle estimations. Consequently, TOP-Nav achieves open-world navigation that the robot can handle terrains or disturbances beyond the distribution of prior knowledge and overcomes constraints imposed by visual conditions. Building upon extensive experiments conducted in both simulation and real-world environments, TOP-Nav demonstrates superior performance in open-world navigation compared to existing methods. Project page at top-nav-legged.github.io.
Keywords: Navigation, Task Planning, Reinforcement Learning
1 Introduction
Figure 1: TOP-Nav achieves open-world navigation in both simulation and the real world. The robot plans an obstacle-free path on terrains with better traversability. For novel terrains, the robot rapidly estimates its traversability based on proprioception experience.
Despite recent advancements in legged locomotion allowing the robots to navigate various terrains based on a simulation-learned strong controller [1, 2, 3, 4, 5, 6, 7, 8, 9], the complexity of currently insimulable real-world factors makes it impossible for the robot to traverse all potential terrains encountered in reality. As a result, integrating the locomotion controller with only an open-loop path planner often restricts legged navigation to limited scenarios [10, 11, 12, 13].
An effective solution to overcome these limitations is equipping the robot with terrain awareness: a traversable path can be planned based on the robot’s preferences on terrains [14, 15] with an appropriate distribution of contact heights and forces [16]. Unlike obstacle estimation, the distinct terrain features are typically encoded as semantic information, where traditional methods collect sufficient data and train segmentation or classification models [17] to learn these features. Nevertheless, compiling an exhaustive catalogue of all conceivable terrains along with their corresponding walking preferences is impractical [18]. Compounding the issue, the dynamic real-world conditions, such as lighting, humidity, and temperature, may introduce inaccuracies in the correspondence between images and walking preferences, especially when relying exclusively on vision in this context [19].
To address the mentioned challenges that relying solely on vision and ignoring motion states in path planning, we complement the vision-only terrain estimator with online corrections derived from motion evaluations. We construct a proprioception advisor to not only convey information about the traversability cost of novel terrains but also alert the robot to unexpected disturbances, such as invisible obstacles.
By integrating the Terrain estimator, Obstacle estimator, and Proprioception advisor, we formulate TOP-Nav: a hierarchical path and motion planning framework navigating a quadruped robot through diverse and challenging terrains proficiently. We evaluate TOP-Nav both in simulation and on a physical robot, with a comparative analysis against existing legged navigation systems. Our main contributions are summarized below:
  • A comprehensive legged navigation framework Integrating multi-modal observations throughout a task and motion planner;
  • A terrain estimator trained from previously collected data to inform the robot of a vision-based terrain traversability;
  • Compensating proprioception to offer online corrections for the vision-based estimation of both terrain traversability and obstacles;
  • Successful implementation and quantitative validations of the proposed framework in both simulation and hardware.
2 RELATED WORK
2.1 Vision and Legged Proprioception Integration
Vision-aided legged navigation has been extensively explored in existing literature [12, 11, 20], with performance heavily dependent on a robust perception module [21]. This reliance makes it difficult and costly to transfer a specific system to different hardware platforms. To provide a comprehensive task observation and reduce the reliance on vision systems, recent researches have introduced proprioception to improve task planning. A majority of these works learn proprioception representations along with visual features in simulation, and then implement the cross-modal features through end-to-end [22], hybrid [23] or decoupled frameworks [24]. Despite the effectiveness demonstrated in these works, the high-dimensional representation space presents challenges for adaptation to novel scenarios and sim-to-real transfer. Alternatively, Fu et al. [25] introduced a hierarchical navigation framework that derives evaluation scores from motion states, yet it overlooks the integration of visual observations. To mitigate these limitations, we propose a novel approach within TOP-Nav by maintaining a series of lightweight cost maps derived from multi-modal observations. This integration achieves a dynamic balance between vision and proprioception. Furthermore, we leverage the learning-based locomotion controller to derive motion evaluations from the value function, offering an efficient solution without additional training.
2.2 Terrain Traversability Estimation
Terrain traversability is determined by factors such as terrain geometry, texture, and physical properties [26]. Estimating these features could be achieved by identifying the semantic class with a predefined static traversability score[27, 17, 28, 14]. These solutions exhibit a notable dependency on large-scale datasets [29] or limited to structured environments like urban scenarios [30, 31]. In off-road navigation, the motion states involved in the dynamic interactions between the robot and the environment provide valuable metrics for assessing terrain traversability [32]. These insights have inspired methods that eliminate the need for manual annotation by autonomously deriving terrain traversability from proprioception through self-supervised learning [19, 33, 34, 35, 36, 37]. Nevertheless, the performance of these studies is contingent upon the quality of the collected datasets [26]. To emphasize the challenges in unconstrained navigation, researchers have proposed various approaches to handling novel observations. For instance, Frey et al. [18] updated the traversability estimation network online with anomalies into consideration. Karnan et al. [38] performs nearest-neighbor search in the proprioception space to align visually novel terrains with existing traversability. Drawing inspiration from those works estimating traversability for novel terrains, we propose a terrain estimator that employs the proprioception advisor as online corrections. Our method diverges from previous approaches primarily in two key aspects: 1) We employ an estimated value function from reinforcement learning to assess terrain traversability, providing a comprehensive evaluation of robot-terrain interactions. 2) We make online corrections on the vision-based estimation without additional training, providing a data-efficient solution for identifying novel terrains.
3 Background
Learning-based legged locomotion controllers have been well developed through reinforcement learning, which is generally achieved by updating the policy ${\pi_{1}}$ within the asymmetric actor-critic training:
${\pi_{1}}\left\{\begin{aligned} \bm{a}_{t}={\pi_{\textrm{actor}}}(\bm{o}_{t}^{p},\bm{o}_{t}^{i},\bm{o}_{t}^{e},\bm{o}_{t}^{h})\\ c_{t}={\pi_{\textrm{critic}}}(\bm{o}_{t}^{p},\bm{o}_{t}^{i},\bm{o}_{t}^{e},\bm{o}_{t}^{h})\end{aligned}\right.,$ (1)
the locomotion policy receives privileged observation ${\bm{o}_{t}^{p}}$, proprioception observation ${\bm{o}_{t}^{i}}$, scanned dots external observation ${\bm{o}_{t}^{e}}$ and historical observation ${\bm{o}_{t}^{h}}$ respectively. ${\pi_{\textrm{actor}}}$ is modelled as a Gaussian policy and infers the optimized actions $\bm{a}_{t}$ to compute the joint positions $\bm{q_{\textrm{des}}}$. $c_{t}$ stands for the estimated value function from $\pi_{\textrm{critic}}$, which is updated through:
$L^{\textrm{critic}}_{t}=(c_{t}-c_{t}^{\textrm{targ}})^{2}\text{ and }c_{t}^{\textrm{targ}}=\sum_{i=t}^{T}\gamma^{i-t}R(s_{i}),$ (2)
${s_{i}}$ denotes the robot state, $R(s_{i})$ represents the rewards accrued at timestep $i$, which commonly includes guiding the robot to track a given velocity commands with stable gaits and attitude. Substantial efforts in reward engineering to formulate $R(s_{i})$ have enabled the value function $c_{t}^{\textrm{targ}}$ to evaluate a comprehensive set of interactions between the robot and its environment. This provides an essential foundation for the proposed proprioception advisor to leverage the estimated $c_{t}$ for motion evaluations within the path planner.
The complete training paradigm will involve a second stage, training a depth encoder and a student network to reproducing ${\bm{o}_{t}^{p}}$ and ${\bm{o}_{t}^{e}}$ from real world accessible observations $\bm{I}_{d}$ (depth image), $\bm{o}_{t}^{i}$ and $\bm{o}_{t}^{h}$. The motion controller will track the target direction $\Delta_{\textrm{yaw}}$ and velocity command $v_{\textrm{lin}}$, the control signal is represented by the desired joint position $\bm{q}_{\textrm{des}}$.
4 Method
4.1 System Overview
TOP-Nav connects a path and motion planner to tackle the task of legged navigation (Fig. 2). A key feature within the path planner (details see the appendix) is the integrated cost map $\bm{M}_{C}$ which offers a comprehensive estimation of external environment that encompasses terrain traversability $\bm{M}_{T}$, obstacle occupancy $\bm{M}_{O}$, proprioception advice $\bm{M}_{P}$ and goal approaching $\bm{M}_{G}$.
Figure 2: TOP-Nav framework: the path planner synthesizes ${\bm{M}_{T},\bm{M}_{O},\bm{M}_{P},\bm{M}_{G}}$ into a combined cost map, from which it computes waypoints based on the overall cost considerations. The motion planner tracks the desired velocity and provide motion evaluations for the proprioception advisor.
4.2 Proprioception Advisor
We design the proprioception advisor to identify motion abnormalities that may arise from unexpected external disturbances. To offer a stable and comprehensive evaluation of the robot-environment interactions, we train the estimated value function (section 3) to synthesize multiple metrics and incorporates historical observations. As demonstarted in Fig. 3(b), the motion evaluation will decline sharply when the robot encounters a locomotion failure caused by transitioning onto challenging terrain. Implementation details for the reward design and normalization are provided in the appendix.
Within the path planner, we first directly utilize the motion evaluations to enhance the robot’s awareness of unforeseen disturbances. This involves estimating invisible collisions ($\bm{M}_{P}$) and integrating them into the combined cost map ($\bm{M}_{C}$). We propose an continuously varied proprioception cost along the y-axis in $\bm{M}_{P}$.
${\bm{M}_{P}}(x,y)=\frac{1-\textrm{Norm}(c_{t})}{e^{k_{P}(\|y_{\textrm{base}}-y\|)}},$ (3)
here $k_{P}$ is hyperparameter. Since a lower motion evaluation indicates a potential challenging terrain or an invisible obstacle in the current direction of the robot, we allocate $c_{t}$ to the centriod column in ${\bm{M}_{P}}$ and decrease it towards the edges. We further discuss the relationship between the proposed motion evaluations, terrain difficulty, and ground truth motion states in the appendix.
Figure 3: (a) The proposed terrain estimator incorporates a visual estimator and online corrections. (b) The motion evaluations respond rapidly when the robot encounters difficult terrains.
4.3 Vision-based Terrain Estimation
The proposed terrain estimator first leverages prior knowledge to map the visual observations to the reference terrain traversability. We apply a perspective transformation on $\bm{I}_{\textrm{rgb}}$ to map the pixels to bird-eye-views $\bm{M}_{\textrm{BEV}}$, which has corresponding coordinates as $\bm{M}_{T}$. We then discretize $\bm{M}_{\textrm{BEV}}$ into patches (Fig. 3(a)) and assign the same difficulty within each patch. The terrain within the prior knowledge $\bm{D}_{\textrm{terrain}}$ is identified through a terrain classification network ${\pi_{\textrm{terrain}}}$.
Considering that such identification falls short when faced with unfamiliar terrains, we calculate the predictive entropy to approximate the identification uncertainty $U$ [39]. With $P_{i}$ denotes the predicted probability of terrain $i$, the predicted explicit terrain names $T$ and corresponding uncertainty $U$ for the current patch can be calculated as:
$U=-\sum_{i}{P_{i}}\log{P_{i}}\text{ and }T=\operatorname*{argmax}_{i}(P_{i}).$ (4)
For safe navigation, terrain traversability signifies the robot’s preference or stability in traversing a particular terrain, which has been explored from either self-supervision [18] or human demonstrations [35]. We leverage experience from these efforts to establish terrain traversability (Fig. 3(a)) based on operator preferences, ensuring alignment with motion evaluations. Detailed discussions are provided in the appendix.
4.4 Online Terrain traversability Corrections
Through the seamless integration of vision and proprioception, we provide online corrections for the vision-based traversability estimation $\bm{M}_{T_{O}}$ without additional training (Fig. 3(a)). This enables the robot with terrain awareness beyond the limitations of the collected data $\bm{D}_{\textrm{terrain}}$.
We address the mechanism by which the robot recalls the traversability of a terrain once it has been seen and traversed within the navigation process. With the robot location $(x_{\textrm{base}},y_{\textrm{base}})$, we record a duration of $1s$ proprioception advice to indicate the terrain traversability at the robot location:
$\bm{T}_{P}(x_{\textrm{base}},y_{\textrm{base}})=1-\textrm{Norm}(\overline{c_{t-k:t}}).$ (5)
Besides $\bm{T}_{P}(x_{\textrm{base}},y_{\textrm{base}})$, we can access a latent feature $\bm{L}(x_{\textrm{base}},y_{\textrm{base}})$ extracted by the classifier ${\pi_{\textrm{terrain}}}$ when the same patch was observed a few steps ago, we record both $\bm{T}_{P}$ and $\bm{L}$ at location $(x_{\textrm{base}},y_{\textrm{base}})$ into an experienced list $\bm{P}_{e}$. Now given a new observation at location $(x_{o},y_{o})$, We can find a traversed and seen patch $(x_{r},y_{r})$ that looks closest to $(x_{o},y_{o})$ based on cosine similarity:
$S_{(o,r)}=\operatorname*{argmax}_{r\in\bm{P}_{e}}\cos{(\bm{L}(x_{r},y_{r}),\bm{L}(x_{o},y_{o}))},$ (6)
Note that we also have the historical traversability $\bm{T}_{P}(x_{r},y_{r})$ infered by the proprioception advisor at $(x_{r},y_{r})$, we can compute a proprioception adapted traversability with the similarity $S_{(o,r)}$ to access the historical proprioception correction ${\bm{M}_{T_{P}}}(x_{o},y_{o})$:
$\bm{M}_{T_{P}}(x_{o},y_{o})=\frac{\bm{T}_{P}(x_{r},y_{r})}{1+e^{-k_{T_{3}}(S_{(o,r)}-S_{0})}},$ (7)
due to the delay in estimation from the proprioception history, we intend for $\bm{M}_{T_{P}}$ to be utilized when encountering novel terrains where $\bm{M}_{T_{O}}$ becomes unreliable. Therefore, we normalize the visual uncertainty $\bm{U}(x_{o},y_{o})$ into confidence $\bm{C}(x_{o},y_{o})$ to adjust the contribution of $\bm{M}_{T_{O}}$ and $\bm{M}_{T_{P}}$:
${\bm{M}_{T}}(x_{o},y_{o})=(\frac{\bm{M}_{T_{O}}-\bm{M}_{T_{P}}}{1+e^{-k_{T_{4}}(C-C_{0})}}+\bm{M}_{T_{P}})(x_{o},y_{o}),$ (8)
here $k_{T_{3}}$, $S_{0}$, $k_{T_{4}}$ and $C_{0}$ are hyperparameters.
By integrating the adapted $\bm{M}_{T}$ as the terrain traversability into the path planner, our system demonstrates rapid adaptability to different terrains and visual conditions.
5 Evaluations
5.1 Experimental Setup
We assess TOP-Nav in challenging navigation environments across both simulations and real-world scenarios, emphasizing the following evaluations: 1) The improvements in navigation performance and locomotion stability achieved by introducing terrain awareness and motion evaluations into the navigation system. 2) The effectiveness of the proposed Proprioception advisor. 3) The effectiveness of the proposed Terrain estimator faced with novel terrains.
Evalutaion Settings Our simulation experiments are conducted within Nvidia Isaac Gym (Fig. 4). We create a grid of ${8\times 8}$ independent navigation cells. Each cell is ${5m\times 5m}$ in size, featuring a robot assigned to a point goal navigation task. The robot and point goal is randomly generated, with a minimum initial distance of ${5m}$ within the cell. The simulation experiments are conducted $25$ times in each of the $64$ navigation cells. For real-world evaluations, we conduct outdoor navigation tasks accross different scenarios involving challenge obstacles and various terrains, with each scenario replicated $5$ times for each method under investigation.
Metrics: We evaluate TOP-Nav with the following metrics: SR (Success Rate): The percentage of successful experiments. We define a success experiment as approaching the point goal into ${0.5m}$ within 20 seconds; TD (Terrain Difficulty): The percentage of average traversed terrain costs, TD provides the ground truth traversed terrain costs within each episode; UT (Unstable Time): The percentage of unstable motion states ($\lvert roll\rvert>0.15\text{ or }\lvert pitch\rvert>0.15$) in each episode; VFT (Velocity Tracking Failure): The average percentage of velocity tracking failures ($\|v_{\textrm{lin}}-v_{\textrm{act}}\|>0.2$) in each episode; AEC (Average Energy Consumption): The average energy consumption ($\tau\dot{q}$) [40] in each successful episode. We calculate the variance across different navigation scenarios to assess the robustness of the proposed method.
Baselines: Beyond ablation study, we compare TOP-Nav against state-of-the-art legged navigation frameworks and segmentation based terrain estimators. VP-Nav [25] integrates vision and proprioception to develop a collision detector and a fall predictor within the navigation pipeline. GA-Nav [17] achieves terrain segmentation relying solely on vision.Sterling [35] learns terrain traversability in a self-supervised manner by assigning traversability to terrains based on those that share similar proprioception representations. However, this approach requires prior training for encountered terrains.
Figure 4: Each navigation cell consists of randomly generated challenging terrains with distinct traverse difficulty, which is marked by the irregularity and complexity of the terrain. The proposed terrain awareness navigation framework plans an optimal path to navigate challenging terrains. The robot demonstrates the capability to recover from unexpected obstacles or irregular terrains with the proprioception advisor.
5.2 Simulation Evaluations
Table 1: Simulation Results with comparison experiments and ablation study
Method SR (%) $\uparrow$ TD(%) $\downarrow$ UT(%) $\downarrow$ VFT(%) $\downarrow$ AEC $\downarrow$
VP-Nav ${65.62\pm 2.4}$ ${47.18\pm 0.8}$ ${36.18\pm 0.5}$ ${26.09\pm 0.4}$ ${101.80\pm 49.6}$
wo/Terrain ${68.00\pm 2.6}$ ${45.83\pm 0.7}$ ${34.92\pm 0.6}$ ${25.06\pm 0.5}$ ${101.70\pm 46.6}$
wo/Proprioception ${62.75\pm 1.6}$ ${\bm{20.47\pm 0.2}}$ ${32.87\pm 0.5}$ ${26.66\pm 0.6}$ ${\bm{88.17\pm 15.3}}$
Obstacle-Only ${52.81\pm 2.3}$ ${47.79\pm 0.8}$ ${41.45\pm 0.8}$ ${39.13\pm 1.0}$ ${96.82\pm 48.0}$
TOP-Nav ${\bm{73.62\pm 1.2}}$ ${22.65\pm 0.2}$ ${\bm{27.21\pm 0.2}}$ ${\bm{18.14\pm 0.2}}$ ${92.08\pm 24.4}$
Improvements with Terrain Awareness: TOP-Nav empowers the robot to select terrains with higher traversability, leading to a significant improvement in the success rate of navigation: TOP-Nav surpasses the VP-Nav baseline by approximately $8\%$ in success rate (SR) and achieve a TD of nearly $20\%$ (Table 1), which is half of the TD achieved in methods without terrain awareness.
Improvements with Proprioception Advisor: The proposed proprioception advisor could recover the robot from locomotion failures such as getting stuck on irregular terrain surfaces or colliding with unseen obstacles during directional changes (Fig. 4). In contrast to methods without proprioception, our approach exhibits an approximate $11\%$ enhancement in SR. Meanwhile, even without terrain awareness, the proposed system outperforms VP-Nav by $2.5\%$, signifying an improvement in our integration of the proprioception advisor compared to existing methods.
Advancements in Locomotion: TOP-Nav achieves the lowest UT and VFT, indicating that selecting simpler terrains contributes to locomotion stability. We evaluate energy consumption with the assumption that when traversing simpler terrain, the robot should exhibit more natural gaits. As a result, TOP-Nav exhibit a $10\%$ reduction in energy consumption compared to methods without terrain awareness.
5.3 Real World Evaluations
Table 2: Real World Evaluation Results on Novel Terrains
Scenerios Unkown Gravel Slippy Tarpaulin
TOP-Nav VP-Nav GA-Nav Obstacle-Only TOP-Nav VP-Nav Sterling
[…]
AEC $\downarrow$ $\bm{57.46}$ 70.3070.30{70.30}70.30 62.5762.57{62.57}62.57 72.6072.60{72.60}72.60 $\bm{83.40}$ 102.0102.0{102.0}102.0 113.8113.8{113.8}113.8
Time(%$\downarrow$ 1 $\bm{62.72}$ 77.9077.90{77.90}77.90 67.4267.42{67.42}67.42 70.6470.64{70.64}70.64 67.6267.62{67.62}67.62 $\bm{63.05}$ 76.5276.52{76.52}76.52
  • 1
    The Time metric measures the proportion of time the robot takes to complete the navigation task.
To evaluate the efficacy of the online corrections for providing the robot with traversability on novel terrains, we conduct experiments using a degraded terrain classifier trained without gravel (Fig 5(A)) and in terrains without any prior information (Fig 5(B)). We demonstrate that the robot can recall previously encountered challenging terrain during the phase of online corrections and plan a waypoint to avoid it. This showcases the fast adaptation ability of TOP-Nav to plan an optimal path when faced with novel terrains.
The quantitative results are provided in Table 2. In the Unkown Gravel experiments, despite GA-Nav possessing complete prior knowledge and correctly identifying the paved road, our online corrections only lead to a $7\%$ increase in UT, mainly due to the forward movement in phase 1. VP-Nav allows the robot to exit the gravel terrain with the safety advisor, but it can not retain this information. In the Slippy Tarpaulin experiments, TOP-Nav demonstrates effectiveness in maintaining stability and improving the success rate by avoiding challenging terrain. In the Slippy Tarpaulin experiments, TOP-Nav demonstarting effectiness in maintaining stability and improving the success rate by avoiding challenging terrain. Among the comparisons, Obstacle-Only keeps moving forward on the slippy surface and exhibit the worst motion performance. We include Sterling in this comparison to demonstrate the importance of online corrections when incorporating proprioception into terrain estimation. Without online corrections, Sterling behaves similarly to Obstacle-Only, unable to adapt to novel terrains and continuing forward. The results against novel terrains highlight the system’s ability to perform open-world navigation.
Figure 5: (A) The terrain classifier does not include high-cost gravel in prior. (B) The robot encounters terrains with no prior knowledge, including a slippery detergent surface. We evaluate the effectiveness of the proposed terrain estimator within TOP-Nav in the experimenrs.
In addition, we conducted a consecutive long-distance experiment (Fig. 5-C) in an off-road scenario with grass and gravel on both sides of the path. The robot is initially given a target direction straight ahead, relying on the terrain and obstacle estimator to keep navigating on the paved road. We observe that in such scenarios, the challenging terrains help guide the robot in changing its moving direction, ensuring that the robot stays on the paved road automatically. More hardware evaluations are provided in the appendix and the supplementary video.
Conclusion
We present TOP-Nav, a legged navigation system that achieves closed-loop integration of visual and proprioception at both task and motion planning levels. Specifically, not only have we incorporated vision-based obstacle and terrain estimation into the navigation framework, but we have also utilized the proprioception advisor to make online corrections on these vision-only estimations. Through the extensive quantitative experiments conducted in both simulation and real-world, we underscore the success of our system in achieving open-world navigation, surpassing limitations posed by terrain, obstacles, or prior knowledge. Compared to existing approaches that couple vision and proprioception or navigate with terrain awareness, our system excels not only in achieving a higher success rate in challenge navigation, but also in demonstrating more stable gaits and lower energy consumption. In the future, we aim to extend the application of the proposed system to longer-distance navigation tasks through precise localization from lidar. Furthermore, we plan to enhance the task planner with data-sufficient models to improve the smoothness and stability.
Appendix
6 Additional Experiments
In this section, we provide additional experiments to justify the effectiveness of the key components in TOP-Nav.
6.1 Evaluations on Vision-Based Terrain Estimator
We conduct additional quantitative hardware experiments on terrains included in the prior datasets (Fig 6), demonstrating superior performance compared to existing legged navigation systems.
Figure 6: We conduct the evaluations in scene 1-3, the green line represents the trajectory of TOP-Nav, while the red line represents Obstalce-Only.
The evaluation results are presented in Table 3. TOP-Nav successfully completes all 15 trials with superior walking stability and minimal energy consumption. These results affirm the successful deployment of the proposed system on real hardware, allowing the robot to select waypoints on terrains with better traversability and thus execute more natural gaits. In contrast, GA-Nav, trained on RUGD [41], demonstrates notable limitations when confronted with open-world navigation. We observe that the improvements of TOP-Nav over GA-Nav mainly stem from: 1) In scene 1,2, GA-Nav fails to identify the slabbed road as a low-cost terrain from the bush and gravel. 2) In scene 3, GA-Nav incorrectly identifies the lower part of the obstacle as a cement floor, which could be addressed by our integration of obstacles and terrain estimation.
Table 3: Evaluation Results on Vision-Based Terrain Estimator
Method SR $\uparrow$ UT(%) $\downarrow$ AEC $\downarrow$ Time(%) $\downarrow$
GA-Nav ${13/15}$ 19.8019.80{19.80}19.80 74.4274.42{74.42}74.42 68.7668.76{68.76}68.76
VP-Nav ${11/15}$ 23.0223.02{23.02}23.02 71.4571.45{71.45}71.45 80.1280.12{80.12}80.12
Obstacle-Only ${12/15}$ 22.7522.75{22.75}22.75 71.0971.09{71.09}71.09 79.7279.72{79.72}79.72
[…]
Nevertheless, GA-Nav surpasses VP-Nav and Obstacle-Only, which navigates without terrain awareness. With Obstacle-Only, the robot gets stuck by the stones or Grassroots, leading to a higher time consumption. While VP-Nav effectively alerts the robot to locomotion failures that could be brought by such terrains, the absence of terrain awareness prevents the robot from successfully navigating out of these challenging terrains, resulting in inferior locomotion performance.
6.2 Evaluations on Proprioception Advisor
In this section, we discuss the relationship between the proposed proprioception advisor and terrain difficulties, along with ground truth reward terms such as velocity tracking and attitude stability.
Figure 7: We conduct a series of straight navigation tasks in simulation. Each robot’s waypoint is set in a straight line, covering terrains of various difficulties $[0.25,0.5,0.75,1]$ with types covering slopes, discrete steps and irregular surfaces.
We conduct a series of straight navigation tasks in simulation, as demonstrated in Fig 7. Each robot’s waypoint is set in a straight line, covering terrains of various difficulties $D_{l}$ and types:
  • Irregular Surface: The height of each point in the terrain is randomly changed within the range of $(0.01+0.04*D_{l},0.07+0.04*D_{l})m$.
  • Discrete Steps: This terrain type consists of $12*D_{l}$ mall steps within a $1m\times 1m$ area, where each small step has dimensions of $(0.1,0.1,0.08)m$.
  • Slope & Holes: Randomly switches between upper slopes and holes, where the maximum height/depth of each slope/hole is set as $1.2m*D_{l}$.
Table 4: Evaluation of the proprioception advisor on different types of terrains.
Method Terrain Slope& Holes Discrete Steps Irregular Surface Flat
Difficulty Ground
Prop Advisor 0.25 $0.7560$ $0.7964$ $0.7001$ $0.8600$
0.50 $0.3316$ $0.3633$ $0.2589$
0.75 $0.1893$ $0.1614$ $0.1058$
1.0 $0.0573$ $0.0837$ $0.0347$
Decoupled Training 0.25 $0.8649$ $0.9836$ $0.9811$ $0.9900$
0.50 $0.3076$ $0.7403$ $0.8141$
0.75 $0.2312$ $0.2323$ $0.4831$
1.0 $0.0770$ $0.3230$ $0.2563$
We conduct $30k$ steps of navigation for each robot and compute the average motion evaluations estimated by the proprioception advisor. The results are shown in Table 4. We observe that as terrain ruggedness, irregularity, and the presence of obstacles increase, there is a corresponding decrease in the motion evaluations. Although this correlation is not strictly linear, the results provide supportive for us to estimate terrain traversability based on the proprioception advisor.
For comparison, we train an independent motion evaluation function $\pi_{\textrm{deq}}$, which is decoupled from the training process of the motion controller. $\pi_{\textrm{deq}}$ is updated using the MSE loss between the ground truth motion states and the network prediction $c_{\textrm{ref}}$. The ground truth walking states include velocity tracking, orientation penalty, and energy consumption:
$c^{\textrm{target}}_{\textrm{ref}}=min(\overline{v_{act}},v_{\textrm{lin}})-(roll^{2}+pitch^{2})-0.001*\tau\dot{q},$ (9)
we train $\pi_{\textrm{deq}}$ for $10k$ iterations, $c^{\textrm{target}}_{\textrm{ref}}$ is normalized using a sigmoid function.
As shown in Table 4, $c_{\textrm{ref}}$ demonstrates a noticeable decline as terrain difficulty increases, yet it does not exhibit a significant advantage compared to the proposed proprioception advisor.
On the other hand, the same terrain may exhibit different traversability for robots with varying locomotion capabilities. Therefore, the evaluation of walking states represents another important metric for assessing terrain traversability. We demonstrate the correlation between the predicted motion evaluations $c_{t}$ and the ground truth motion states $c^{\textrm{target}}_{\textrm{ref}}$ in Fig 8.
Figure 8: The estimated $c_{t}$ and the actual motion states of the robot has a strong linear correlation.
The perason correlation coefficient is $0.914$, indicating a strong linear correlation between the estimated $c_{t}$ and the actual motion states of the robot.
Building upon the discussion above, we validate that constructing the proprioception advisor with critic output enables efficient motion evaluation, including information on both the terrain difficulty and walking states of the robot. Moreover, the proposed advisor can be implemented without the need for additional training or sensors, making it a more appropriate approach to provide motion evaluations for robot path planning compared to existing methods.
6.3 Different Velocities
We conducted quantitative experiments in simulation with different velocities ($0.25,0.5,0.75,1.0$) (m/s), and the results demonstrate that proper velocity design does affects the performance of the navigation system. As demonstrated in Table 5, a slower velocity command (0.25) prevents the robot from reaching the goal on time, resulting in a lower success rate. Moreover, it can be concluded that lower the speed would help decrease the energy consumption.
Table 5: Simulation Results with velocity comparisons
Velocity (m/s) SR (%) $\uparrow$ TD(%) $\downarrow$ UT(%) $\downarrow$ VFT(%) $\downarrow$ AEC $\downarrow$
$0.25$ ${49.69\pm 3.43}$ ${26.22\pm 0.33}$ ${26.56\pm 0.41}$ ${25.95\pm 0.45}$ ${65.55\pm 352.19}$
0.5 ${73.62\pm 1.21}$ ${22.65\pm 0.19}$ ${27.21\pm 0.24}$ ${18.14\pm 0.16}$ ${92.08\pm 24.39}$
$0.75$ ${73.91\pm 2.46}$ ${22.22\pm 0.26}$ ${22.53\pm 0.37}$ ${16.79\pm 0.28}$ ${80.32\pm 46.82}$
$1.0$ ${72.03\pm 2.80}$ ${21.65\pm 0.23}$ ${41.92\pm 0.49}$ ${21.30\pm 0.44}$ ${123.53\pm 39.61}$
On the other hand, setting a too large velocity command will introduce locomotion instability, as indicated by the results with a velocity of $1m/s$. However, conducting a quantitative analysis of the influence of speed is challenging in complex terrains. For instance, in slopes, robots with higher velocities may navigate more effectively, whereas terrains with discrete steps could lead to locomotion failures at faster speeds. Consequently, experiments with velocity commands of $1m/s$ still achieve a high success rate.
The performances vary between $0.5m/s$ and $0.75m/s$ is not significant; therefore, we consider this range to be an appropriate commanded velocity range for deployment. In real-world experiments, we demonstrate the robot’s velocity set at $0.75m/s$ in the supplementary video.
6.4 Invisible Obstacles
Figure 9: The robot avoids invisble obstacles with the proprioception advisor.
Metrics VP-Nav Obstacle-Only TOP-Nav
[…]
Table 6: Evaluation results on the Invisible Obstacles.
The proposed proprioception advisor provides close-loop feedback to alert the robot to invisible obstacles. We assess this capability in Fig. 9, where a glass wall is located along the planned path. In phase 1, the obstacle estimator is oblivious to the presence of the glass wall, and the robot continues moving forward. In phase 2, the robot collides with the glass wall, causing a rapid increase in ${\bm{M}_{P}}$ along the direction of movement. As a result, the robot adjusts its waypoint to steer clear of the high-cost area, successfully avoiding the glass wall. The quantitative results are provided in Table 6.
7 Implementation Details
7.1 Task Configuration
To elucidate the task configuration, the robot is given a point goal with its location $\bm{p}_{\textrm{goal}}$ and has to approach the goal within 10 seconds. The location of the robot is denoted as $(\bm{p}_{\textrm{base}},r_{\textrm{base}})$. For external observation, we utilize a bio-channel perception module for both the task and motion planner. This module includes depth ($\bm{I}_{d}$) and RGB ($\bm{I}_{\textrm{rgb}}$) channels. The path planner computes the desired velocity command $(v_{\textrm{lin}},\Delta_{\textrm{yaw}})$ for the robot, which is tracked by the motion planner along with the appropraite joint position signal.
7.2 Path Planner Details
TOP-Nav integrates the combined cost map $\bm{M}_{C}$, offering a comprehensive estimation that encompasses terrain traversability $\bm{M}_{T}$, obstacles $\bm{M}_{O}$, proprioception advice $\bm{M}_{P}$ and the goal approaching map $\bm{M}_{G}$. This integration addresses both visible and unexpected external disturbances against the robot.
To keep a balance between navigation efficiency and safety, we formulate the combination of $\bm{M}_{C}$ with dynamic scaling factors as follows:
$\bm{M}_{C}=\bm{M}_{P}+\bm{M}_{O}+{\alpha_{T}}\bm{M}_{T}+{\alpha_{G}}\bm{M}_{G},$ (10)
$\alpha_{T}=\frac{k_{T_{1}}}{1+e^{-k_{T_{2}}(d-d_{0})}}\text{, }\alpha_{G}=\frac{k_{G_{1}}}{1+e^{-k_{G_{2}}(t-t_{0})}}+k_{G_{0}},$ (11)
here $\bm{M}_{i}$ are 2-d matrix with a spatial resolution of $0.15m$, $d_{0},t_{0},k_{T},k_{G}$ are hyperparameters, ${t}$ denotes current time consuming and ${d}$ denotes the distance from the point goal at the current step. We accord the highest priority to $\bm{M}_{P}$ and $\bm{M}_{O}$ since they represent non-traversable locations where the robot cannot pass through. The terrain traversability scale ${\alpha_{T}}$ decreases as the robot approaches the target. This design is made considering that, as the robot nears the target (${d}$ decreases), taking a detour to avoid a challenging yet traversable terrain would be an inefficient behaviour. Conversely, the goal approaching scale ${\alpha_{G}}$ increases with the duration ${t}$ of the task.
With the integrated map $\bm{M}_{C}$, we select the optimal waypoint $\bm{p}_{\textrm{way}}$ based on the lowest combined cost.
$\bm{p}_{\textrm{way}}=\operatorname*{argmin}_{\bm{p}\in\bm{E}}\frac{1}{\|\bm{p}_{\textrm{base}}-\bm{p}\|}\sum_{(x,y)=\bm{p}_{\textrm{base}}}^{{(x,y)=\bm{p}}}\bm{M}_{C}(x,y),$ (12)
here $\bm{E}$ denotes the set of points on the edge of $\bm{M}_{C}$. For each point $\bm{p}$ in $\bm{E}$, we calculate the average combined cost along the path from the robot location $\bm{p}_{\textrm{base}}$ to the edge point $\bm{p}$. The optimal waypoint $\bm{p}_{\textrm{way}}$ is then chosen as the point with the lowest path cost.
After determining the optimized local target $\bm{p}_{\textrm{way}}$, the target direction is calculated based on the relative position $\Delta_{\textrm{yaw}}=\arctan(\bm{p}_{\textrm{way}}-\bm{p}_{\textrm{base}})-r_{\textrm{base}}$, here $r_{\textrm{base}}$ denotes the yaw direction of the robot base. The linear velocity is constrained to mitigate the impact of high angular variation with $v_{\textrm{lin}}=v_{0}e^{-k\Delta_{\textrm{yaw}}}$, thereby preventing abrupt and substantial turning at high speeds.
In simulation, $\bm{M}_{C}$ is configured with dimensions $(1.2m,1.5m)$ centered around the robot‘s location at ${\bm{p}_{\textrm{base}}^{\textrm{map}}=(0.45m,0.75m)}$. This design prioritizes a light-weight, real-time updated path planner and maintains consistency with the height map used in locomotion. The environments include randomly generated obstacles and terrains. Each cell is equipped with at least one wall obstacle, two column obstacles and the remaining space is divided into ${1m\times 1m}$ sections. For terrain traversability assignments, we uniformly partition the obstacle-free space into difficulty levels [0, 0.25, 0.5, 0.75, 1] and generate terrain with various heights of steps and the intensity of irregular terrain corresponding to the difficulty levels. This difficulty serves as ground-truth walking preferences in simulation.
To obtain a more comprehensive observation in the real world, we configure the cost map $\bm{M}_{C}$ with dimensions of $(3m,3m)$. The location of robot is at ${\bm{p}_{\textrm{base}}^{\textrm{map}}=(0m,1.5m)}$ in real world $\bm{M}_{C}$. The criteria for success include reaching the goal within the specified time constraints, consistent with the simulation setting.
Obstacle Estimation and Localization: In simulation, the robot has access to the ground truth location. For each point $\bm{p}=(x,y)$, $\bm{M}_{G}(x,y)=\|\bm{p}_{\textrm{goal}}-\bm{p}\|$. In real world, we set a target direction $r_{\textrm{goal}}$ to compute the goal map $\bm{M}_{G}(x,y)=\|r_{\textrm{goal}}-\arctan(\bm{p}-\bm{p}_{\textrm{base}})\|$. We construct the obstacle estimation based on the depth channel $\bm{I}_{d}$. The perceived obstacles are converted into point clouds, and for each point in $\bm{M}_{O}$, we compute the signed distance $\bm{M}_{\textrm{SDF}}$ to the closest obstacles within distance $d_{\textrm{max}}$, the cost of obstacles therefore can be computed as:
$\bm{M}_{O}(x,y)=\frac{\max(0,d_{\textrm{max}}-\bm{M}_{\textrm{SDF}}(x,y))}{d_{\textrm{max}}}.$ (13)
For reference, we provide the default values of the hyperparameters used in the path planner in Table 7. Parameters $d_{0},t_{0}$ adjust the weights of $\bm{M}_{T}$ and $\bm{M}_{G}$ as the robot approaches the target and as time progresses. $k_{G_{0}}$ keeps a minimum weight for the robot approaching the target. The listed values remain consistent in simulation; however, considering the various conditions in the real world, we make slight changes to these values in different real-world experiments for better validation of the contribution of this work. For instance, in outdoor navigation experiments, we assign a higher value to $k_{T_{1}}$ to expand the scale of the terrain estimator, which is effective for evaluating the accuracy of the proposed terrain estimator.
The default commanded velcoity $v_{0}$ is set to $0.5m/s$ in both simulation and the real world. The maximum distance $d_{\textrm{max}}$ we considered for computing the obstacle map $\bm{M}_{O}$ is $0.3m$.
Table 7: Hyperparameters in the path planner
Parameter $d_{0}$ $t_{0}$ $k_{T_{1}}$ $k_{T_{2}}$ $k_{G_{0}}$ $k_{G_{1}}$ $k_{G_{2}}$
Default Value $0.5$ $0.5$ $1.0$ $2.0$ $0.1$ $0.4$ $-10$
7.3 Motion Planner Details
The motion planner is implemented following [3, 42]. Both the actor and critic networks within the locomotion policy ${\pi}$ have hidden layer sizes of [512, 256, 128]. The proprioception observation $\bm{o}_{t}^{p}$ includes angular velocity (3), Orientation (2), velocity commands (3), joint positions (12), joint velocities (12), and the last action (12). We store the last 10 steps of $\bm{o}_{t}^{p}$ into the history observation $\bm{o}_{t}^{h}$. The depth encoder learns the exteroception latent features from onboard observation $\bm{I}_{d}$ using a conv-GRU structure. ${\pi}$ is updated using PPO [43] for 20K iterations, the batch size is 160000 divided into 4 mini-batches. The depth encoder is optimized for 5k iterations. The rewards obtained at each step are the sum of the reward functions listed in Table 8.
Table 8: Reward Functions
Target Velocity Tracking1 $min(\overline{v_{act}},v_{\textrm{lin}})$
Target Direction Tracking $e^{-\mid\Delta_{\textrm{yaw}}\mid}$
Orientation Penalty $-(roll^{2}+pitch^{2})$
Hip Joint Position Penalty ${-\|q_{hip}\|}^{2}$
z-direction velocity Penalty ${-\mid v_{z}\mid}^{2}$
Collision Penalty2 $-\sum_{i\in(calf,thigh)}({F_{c}^{i}\geq 0.1N})$
Torques Variation Penalty ${-\|\tau_{t}-\tau_{t-1}\|}$
  • 1
    $\overline{v_{act}}$ is the projection component of $v_{act}$ in the target direction.
  • 2
    $F_{c}^{i}$ is the contact force of calf and thigh indices.
7.4 Proprioception Advisor Details
The default $\pi_{\textrm{critic}}$ requires privileged observations, we train another value function estimation network using observable motion states for the proprioception advisor to be deployed on hardware. The estimated motion evaluation $c_{t}$ is normalized using a sigmoid function. The proprioception advisor is trained with the locomotion policy ${\pi}$, with an exact same network architechture as $\pi_{\textrm{critic}}$. The input only includes proprioception observation $\bm{o}_{t}^{p}$ and historical observation $\bm{o}_{t}^{h}$, updated through the MSE loss with $c_{t}^{\textrm{targ}}$. The estimated value $c_{t}$ is normalized through sigmoid function:
$Norm(c_{t})=\frac{1}{1+e^{2*(2.2-c_{t})}}$ (14)
In practice, and we set a minimum threshold $c_{th}$ for the normalized critic value be considered in the path planner, in simulation $c_{th}=0.8$, while in the real world $c_{th}=0.5$, $Norm(c_{t})$ larger than $c_{th}$ will be set to $1$.
$\bm{M}_{P}$ is calculated with $k_{P}=0.3$, we demonstrate an example of $\bm{M}_{P}$ after the robot collides with a wall (with vision and obstacle detection disabled in this trial) in Fig 10. In this scenario, the normalized $c_{t}$ returns $0.02$, $\bm{M}_{P}$ has a width of $0.75m$ along the moving direction of the robot with costs greater than 0.5.
Figure 10: We demonstrate an example of $\bm{M}_{P}$ when the robot collides with an obstacle.
Generally, the proprioception advisor engages when locomotion states are compromised. To enable recovery from being obstructed by the unexpected terrains, we have designed the following recovery strategy in simulation: when $Norm(c_{t})$ falls below a specific threshold (0.5), we initially assign a backward velocity command of 0.5 m/s for 0.5 seconds. Next, to address potential terrain-related constraints that may prevent the robot from reaching planned waypoints at its default velocity, we reduce the frequency of the path planner to 0.5 Hz until the robot reaches the waypoint or makes a directional change of 0.3 radians.
7.5 Terrain Estimator Details
For prior data collection of the terrain estimator, we teleoperate the robot to walk for 5 minutes on each of the concerned terrains. We capture the motion evaluations $c_{t}$ as well as the first-view observations during these demonstrations. We convert the RGB observations into BEV maps and dividing them into patches. The collected data is labelled based on the corresponding terrain from which it was acquired. The classification network is implemented using the MobileNet backbone and trained on a Nvidia GTX 2080 Ti for 6 hours. The proposed pipeline provides a light-weight inference model of 8 MB to be deployed for onboard computation. During deployment, to approximate the uncertainty $U$ with predictive entropy, We perform $K=8$ stochastic forward inferences with different data augmentations as Monte Carlo simulation, We estimate the expectation as the average of the predicted probability $P_{i}=\sum_{k}P_{i}^{k}/K$.
Based on the collected motion evaluations, we compute the average $\bm{T}_{P}$ observed during walks on each terrain. These walking experiences provide reference values for assigning the terrain costs defined by the operator. Table 9 presents the validation accuracy of the terrain classifier, along with the reference terrain cost and the average collected $\bm{T}_{P}$ for each terrain. Since the terrains encountered in the real world did not exhibit significant differences in motion evaluations, we did not strictly set the reference cost based on linear correlation.
Table 9: Terrain Cost and Classify Accuracy
Name Bush Brick Snow Gravel Slab Cement Paved Grass
[…]
$\bm{T}_{P}$ 0.481 0.523 0.445 0.507 0.581 0.580 0.580 0.572
Cost 0.7 0.6 0.9 0.7 0.0 0.0 0.0 0.2
The hyperparameters used for online corrections are fine-tuned based on the distribution of the learned latent features $\bm{L}$ from the pre-collected data. Figure 11-(b) depicts a t-SNE visualization of the features inferred by the terrain classifier, involving data both within and outside of $\bm{D}_{\textrm{terrain}}$. We observe that $\pi_{\textrm{terrain}}$ can learn distinctive features for specific terrains, showcasing unique clustering in the latent space, even for previously unseen terrain types. The average cosine similarity $S_{(o,r)}$ within each terrain class is $0.95$, we set the middle point $S_{0}$ to be $0.85$ and $k_{T_{3}}$ to be $15$. This allows the terrain cost derived from historical experience to decrease rapidly when the similarity falls below $0.8$.
Figure 11: (a) We illustrate the distribution of samples across different confidence levels, with the training of the terrain classifier, the confidence values for known terrains predominantly clustered beyond 0.9, while samples of novel terrains remained dispersed across the confidence axis. (b)T-SNE visualization of the features inferred by the terrain classifier.
To demonstrate the effectiveness of computing confidence in deciding whether the observed terrains is outside of $\bm{D}_{\textrm{terrain}}$, we illustrate the distribution of samples across different confidence levels in Fig 11-(a). After 200 steps of training, the confidence values for known terrains predominantly clustered beyond 0.9, while samples of novel terrains remained dispersed across the confidence axis. The confidence is used to adjust the weight of online correct ions in terrain estimation. We set $k_{T_{4}}=20$ and $C_{0}=0.9$ in the experiments.
7.6 Hardware Details
We implement TOP-Nav on the Unitree-Go1 and Unitree-Go2 quadruped robot, with a weight of approximately $12kg$ and dimensions of $645mm\times 280mm$. The robot is equipped with 12 brushless motors, each capable of producing a torque of $35.5Nm$. The perception module is equipped with a RealSense D435i camera, offering simultaneous depth and RGB channels. All computations are processed onboard with an NVIDIA Jetson NX asynchronously, i.e. the motion planner operates at a fixed frequency of $50Hz$ and receives the latest depth latent $\bm{\ell}$ inference from the perception module. The path planner operates at a frequency of ${3Hz}$.